OpenStudy (anonymous):

There are 12 men at a dance. (a) In how many ways can eight of them be selected to form a cleanup crew? (b) How many ways are there to pair off eight women at the dance with eight of these 12 men?

OpenStudy (anonymous):

(a)no. of ways=12c8=12c(12-8)=12c4=12*11*10*9/(4*3*2*1)=11*5*9=495 (b)no. of ways 8 men can be selected to form pairs with women=12c8=495 8 women can be selected in ways=8! total no of ways=495*8!